# Multilayer network science: theory, methods, and applications

### v1

@@ III.10 Network neuroscience
Neuroscience is fundamentally centred on the study of the brain as a vast, interconnected system of neurons and brain regions. This field encompasses not only the structural and functional organization of the brain but also the dynamic processes that underpin cognition, behaviour, and neurological health. Networks provide a natural framework for capturing these complexities by representing the brain as a system of nodes and edges. Over the past decade, network neuroscience bassett2017network; barabasi2023neuroscience has emerged as one of the most exciting and influential fields, offering new insights into the brain’s organization and its intricate functions.
@@
Brain networks bullmore2009complex are mathematically modelled systems where nodes represent either neurons or specific brain regions, while edges capture various forms of connectivity. These connections may represent structural connectivity, which reflects the physical pathways linking brain regions and is often derived from diffusion MRI or diffusion tensor imaging. Alternatively, they may capture functional connectivity, which describes statistical relationships in neural activity between regions, as observed through functional MRI, electroencephalography, or magnetoencephalography. A third type, effective connectivity, goes a step further by modelling causal interactions and directional relationships between regions, often using techniques like Granger causality or dynamic causal modelling nicosia2013. Together, these distinct forms of connectivity provide a rich and multifaceted view of the brain’s organization and activity.
@@
Building on this foundation, multilayer networks offer a sophisticated extension for exploring brain connectivity and dynamics de2017multilayer. These models expand the traditional network framework by incorporating multiple interconnected layers of nodes and edges, each representing a different dynamic or scale of analysis betzel2017multi. Within an individual, for example, these layers can correspond to distinct types of connectivity, such as structural and functional networks, with interlayer links capturing relationships between them vaiana2018multilayer. Similarly, temporal transitions in dynamic functional connectivity can be represented as changes within and across layers or states sporns2021dynamic. Beyond individuals, multilayer networks also enable the comparison of connectivity patterns across populations by modelling interlayer relationships that represent similarities between individuals or groups wilson2020analysis; simas2015. This versatility allows multilayer networks to bridge gaps between data types and scales, revealing insights that would be inaccessible using single-layer approaches mandke2018comparing.
@@
The ability to integrate and analyse these diverse sources of data has led to important studies in areas such as human learning bassett2017learning, individual differences betzel2019community, the diagnosis and understanding of neurological and mental disorders stam2014modern, among others like the relation between social environment and variation in brain network organization merritt2024stability.
@@
Additionally, even though traditional network models focus mainly on pairwise relationships, higher-order interactions can be pulled into the multilayer network paradigm by considering multiple layers describing different orders of interaction. This approach has contributed to new advances in the understanding of human brain evolution and cognition luppi2022synergistic, and the relation between brain activity and behaviour santoro2024higher.
@@
In summary, the multilayer networks framework represents a frontier in network neuroscience, enabling researchers to integrate diverse data types and explore the brain’s multi-dimensional nature. Continued methodological advances will undoubtedly unlock clearer insights into brain function and dysfunction, fostering progress in both neuroscience and clinical applications.
@@ IV Data & Tools
Given the growing interest in network science, a wide array of resources has become available. Yet, despite substantial progress, challenges remain in terms of data availability, standardization, and tool interoperability. This section highlights the most relevant public data repositories and software libraries that support advanced network analysis and visualization.
@@ IV.1 Data
Network data comes from a wide range of disciplines such as biology, sociology, economics, neuroscience, and more. While numerous repositories exist for general-purpose network datasets, multilayer and temporal network datasets are comparatively less common and often tailored to specific domains. Furthermore, the lack of standardized formats can hinder reproducibility and comparative studies across methods.
@@
Netzschleuder 163735 Netzschleuder
Network Repository 5664 nr
KONECT 1326 konect
ICON 704 icon
SNAP 127 snapnets
UCI 35 DuBois:2008
@@
While many general-purpose datasets are readily accessible and well-documented, multilayer and higher-order data remain underrepresented. A key issue is the lack of standardized file formats to represent these complex structures, which can complicate interoperability between tools and comparisons across studies.
@@ IV.2 Software Tools
The development of dedicated software libraries has significantly accelerated the field of network science. A wide variety of tools exist for analysis, simulation, and visualization of networks. Table 2 summarizes major software libraries, including their main focus and programming language.
However, no single tool dominates the field. Even for general-purpose networks, researchers choose between different libraries based on project requirements. For example, while networkX is widely used for its ease of use and versatility, alternatives such as igraph or graph-tool offer superior performance on large networks due to optimized C/C++ backends. Thus, selecting the appropriate tool often involves trade-offs between ease of use, scalability, feature set, and visualization capabilities.
@@
Analysis
DyNetX Dynamic networks Python dynetx
HAT Higher-order networks Python, MATLAB Pickard2023Jun
graph-tool General-purpose networks Python peixoto_graphtool_2014
HGX Higher-order networks Python lotito2023hypergraphx
HyperNetX Higher-order networks Python hypernetx
iGraph General-purpose networks Python, R, C igraph
MultilayerGraphs.jl Mutilayer networks Julia MultilayerGraphs
multinet Mutilayer networks R multinetR
networkX General-purpose networks Python networkx
NDlib General-purpose networks Python Rossetti2018Feb
pathpy Higher-order temporal networks Python Hackl2021Apr
pymnet Mutilayer networks Python pymnet
pyunicorn General-purpose networks Python Donges2015Nov
Reticula Higher-order temporal networks Python badie2023reticula
SimpleHypergraphs.jl Higher-order networks Julia Spagnuolo2020Mar
straph Temporal networks Python Latapy2018Dec
tacoma Temporal networks Julia tacoma
TGLib Temporal networks Python, C++ Oettershagen
tnetwork Temporal networks Python tnetwork
XGI Higher-order networks Python Landry_XGI_2023
Visualization
Cytoscape General-purpose networks GUI shannon2003cytoscape
Gephi General-purpose networks GUI bastian2009gephi
ggraph General-purpose networks R ggraph
Graphia General-purpose networks GUI Freeman2022Jul
MultiNet Multivariate networks Web-based 2023_nsf_multinet
multiNetX Multilayer networks Python multinetX
muxViz Multilayer networks R, Web-based muxviz
Netminer General-purpose networks GUI Ghim2014
Netwulf General-purpose networks Python, Web-based Aslak2019Oct
NodeXL General-purpose networks Excel nodexl
Pajek General-purpose networks GUI Batagelj2004
py3Plex Mutilayer networks Python Skrlj2019
Sigma.js General-purpose networks JavaScript sigmajs
Tulip General-purpose networks GUI Auber2012Jan
Table 2: Libraries and tools for network science analysis and visualization, with support for multilayer, temporal, and higher-order networks. “Main focus” refers to the area where the software is primarily applied; however, some tools can support multiple network types.
@@ V Outlook
The past decade has shown that multilayer networks are not merely a convenient generalization of graph theory but a fundamentally new modelling language. The field has now reached a stage where the central question is no longer how to adapt single-layer tools, but how to identify the truly multilayer mechanisms that cannot be reduced away. Several theoretical directions emerge from this perspective. One concerns the search for minimal descriptions: determining when layers encode irreducible information and when they can be compressed. Another direction concerns the interplay between multilayer topology and cross-layer dynamical coupling, where instabilities, hybrid phase transitions, metastability, or coexistence of dynamical regimes can arise due to the multiplexity nature of the system. Understanding the general principles that govern these phenomena remains a major open challenge. At the same time, the field’s inherently multidisciplinary nature complicates the picture: multilayer approaches are used in physics, biology, economics, computer science, and the social sciences, often with different assumptions, terminologies, and modelling priorities. This diversity makes communication across communities difficult, yet it also underscores the value of multilayer networks as a common representational language capable of unifying disparate viewpoints.
At the methodological frontier, several additional opportunities demand attention. The first involves integrating multilayer representations with temporal and higher-order structures. Many real systems combine heterogeneous interaction types with time-varying group interactions, but current modelling frameworks still tend to treat these features separately. Developing unified theoretical tools capable of capturing polyadic interactions, memory effects, and cross-layer dependencies simultaneously is an important next step. A second direction concerns inference, where the rise of generative models, statistical reconstruction techniques, and machine-learning-based embeddings is beginning to provide new ways to infer hidden layers and mesoscale structures. Yet, understanding the limits of inference remains unresolved. A third challenge lies in controllability and optimization: while single-layer structural controllability is now well understood, its multilayer analogue is largely unexplored, especially when layers interact through non-linear or time-dependent couplings. Finally, the rapidly growing integration between network science and artificial intelligence opens a promising but underdeveloped path. Deep learning architectures are themselves multilayered systems whose internal dynamics could benefit from tools developed in multilayer network science, potentially contributing to interpretability and robustness in AI.
@@
Looking ahead, progress will depend critically on the availability of high-resolution multilayer data, standardized formats, and interoperable tools, but also on conceptual clarity about what it means for a system to be truly multilayer. If these developments continue in parallel, multilayer network science is poised to move from descriptive analyses to predictive, actionable models capable of guiding interventions, optimizing coupled infrastructures, and informing policy in domains where interdependence is no longer optional to ignore. As the complexity of modern systems continues to grow, the multilayer perspective will remain essential for understanding the intertwined processes that characterize the natural and engineered world.
@@ Acknowledgements.
All authors acknowledge the support of the AccelNet-MultiNet program, a project of the National Science Foundation (Award #1927425 and #1927418). A.S.T. acknowledges support by FCT – Fundação para a Ciência e Tecnologia – through the LASIGE Research Unit, ref. UID/408/2025. AA acknowledges support from the grant RYC2021‐033226‐I funded by MCIN/AEI/10.13039/501100011033 and the European Union NextGenerationEU/PRTR. $A.$ and Y.M. were partially supported by the Government of Aragon, Spain, and ERDF ”A way of making Europe” through grant E36-23R (FENOL), and by Grant No. PID2023-149409NB-I00 from Ministerio de Ciencia, Innovación y Universidades, Agencia Española de Investigación (MICIU/AEI/10.13039/501100011033) and ERDF ”A way of making Europe”.

### v2

@@ III.10 Network neuroscience
Neuroscience is fundamentally centred on the study of the brain as a vast, interconnected system of neurons and brain regions. This field encompasses not only the structural and functional organisation of the brain but also the dynamic processes that underpin cognition, behaviour, and neurological health. Networks provide a natural framework for capturing these complexities by representing the brain as a system of nodes and edges. Over the past decade, network neuroscience Bassett and Sporns (2017); Barabási et al. (2023) has emerged as one of the most exciting and influential fields, offering new insights into the brain’s organisation and its intricate functions.
@@
Brain networks Bullmore and Sporns (2009) are mathematically modelled systems where nodes represent either neurons or specific brain regions, while edges capture various forms of connectivity. These connections may represent structural connectivity, which reflects the physical pathways linking brain regions and is often derived from diffusion MRI or diffusion tensor imaging. Alternatively, they may capture functional connectivity, which describes statistical relationships in neural activity between regions, as observed through functional MRI, electroencephalography, or magnetoencephalography. A third type, effective connectivity, goes a step further by modelling causal interactions and directional relationships between regions, often using techniques like Granger causality or dynamic causal modelling Nicosia et al. (2013b). Together, these distinct forms of connectivity provide a rich and multifaceted view of the brain’s organisation and activity.
@@
Building on this foundation, multilayer networks offer a sophisticated extension for exploring brain connectivity and dynamics De Domenico (2017); Presigny and De Vico Fallani (2022). These models expand the traditional network framework by incorporating multiple interconnected layers of nodes and edges, each representing a different dynamic or scale of analysis Betzel and Bassett (2017). Within an individual, for example, these layers can correspond to distinct types of connectivity, such as structural and functional networks, with interlayer links capturing relationships between them Vaiana and Muldoon (2020). Similarly, temporal transitions in dynamic functional connectivity can be represented as changes within and across layers or states Sporns et al. (2021). Beyond individuals, multilayer networks also enable the comparison of connectivity patterns across populations by modelling interlayer relationships that represent similarities between individuals or groups Wilson et al. (2020); Simas et al. (2015). This versatility allows multilayer networks to bridge gaps between data types and scales, revealing insights that would be inaccessible using single-layer approaches Battiston et al. (2018); Mandke et al. (2018); Presigny et al. (2024).
@@
The ability to integrate and analyse these diverse sources of data has led to important studies in areas such as human learning Bassett and Mattar (2017), individual differences Betzel et al. (2019), the diagnosis and understanding of neurological and mental disorders Stam (2014), among others like the relation between social environment and variation in brain network organisation Merritt et al. (2024).
@@
Additionally, even though traditional network models focus mainly on pairwise relationships, higher-order interactions can be pulled into the multilayer network paradigm. For instance, one could consider multiple layers describing different orders of interaction, an approach that has contributed to new advances in the understanding of human brain evolution and cognition Luppi et al. (2022), and the relation between brain activity and behaviour Santoro et al. (2024). In addition, higher-order structures can help uncover the relationship between structure and function in the brain Battiston et al. (2017).
@@
In summary, the multilayer networks framework represents a frontier in network neuroscience, enabling researchers to integrate diverse data types and explore the brain’s multidimensional nature. Continued methodological advances will undoubtedly unlock clearer insights into brain function and dysfunction, fostering progress in both neuroscience and clinical applications.
@@ IV Data & Tools
Given the growing interest in network science, a wide array of resources has become available. Yet, despite substantial progress, challenges remain in terms of data availability, standardisation, and tool interoperability. This section highlights the most relevant public data repositories and software libraries that support advanced network analysis and visualisation.
@@ IV.1 Data
Network data comes from a wide range of disciplines such as biology, sociology, economics, neuroscience, and more. While numerous repositories exist for general-purpose network datasets, multilayer and temporal network datasets are comparatively less common and often tailored to specific domains. Furthermore, the lack of standardised formats can hinder reproducibility and comparative studies across methods.
@@
Netzschleuder 163735 Peixoto (2020)
Network Repository 5664 Rossi and Ahmed (2015)
KONECT 1326 Kunegis (2013)
ICON 704 174
SNAP 127 Leskovec and Krevl (2014)
UCI 35 DuBois (2008)
@@
While many general-purpose datasets are readily accessible and well-documented, multilayer and higher-order data remain underrepresented. A key issue is the lack of standardised file formats to represent these complex structures, which can complicate interoperability between tools and comparisons across studies.
@@ IV.2 Software Tools
The development of dedicated software libraries has significantly accelerated the field of network science. A wide variety of tools exist for analysis, simulation, and visualisation of networks. Table 2 summarizes major software libraries, including their main focus and programming language. It is important to note that this overview focuses on general-purpose tools for the structural analysis and visualisation of complex networks. Tools primarily focused on specific functional processes or computational workflows, such as those for graph neural networks PyTorch Foundation (2025) or epidemic simulations Czuba et al. (2024); Gozzi et al. (2025); Miller and Ting (2019), are not included.
However, no single tool dominates the field. Even for general-purpose networks, researchers choose between different libraries based on project requirements. For example, while networkX is widely used for its ease of use and versatility, alternatives such as igraph or graph-tool offer superior performance on large networks due to optimised C/C++ backends. Thus, selecting the appropriate tool often involves trade-offs between ease of use, scalability, feature set, and visualisation capabilities.
@@
Analysis DyNetX Dynamic networks Python Rossetti (2020)
HAT Higher-order networks Python, MATLAB Pickard et al. (2023)
graph-tool General-purpose networks Python Peixoto (2014)
HGX Higher-order networks Python Lotito et al. (2023)
HyperNetX Higher-order networks Python Praggastis et al. (2024)
iGraph General-purpose networks Python, R, C Csárdi and Nepusz (2006)
MultilayerGraphs.jl Multilayer networks Julia Moroni and Monticone (2022)
multinet Multilayer networks R Magnani et al. (2021b)
NetworKit Large-scale networks Python Research Group Modeling and Analysis of Complex Systems (2025)
networkX General-purpose networks Python Hagberg et al. (2008)
NDlib General-purpose networks Python Rossetti et al. (2018)
pathpy Higher-order temporal networks Python Hackl et al. (2021)
pymnet Multilayer networks Python Nurmi et al. (2024)
pyunicorn General-purpose networks Python Donges et al. (2015)
Reticula Higher-order temporal networks Python Badie-Modiri and Kivelä (2023)
SimpleHypergraphs.jl Higher-order networks Julia Spagnuolo et al. (2020)
straph Temporal networks Python Latapy et al. (2018)
tacoma Temporal networks Julia Maier and Schlosser
TGLib Temporal networks Python, C++ Oettershagen and Mutzel (2022)
tnetwork Temporal networks Python Remy and Jorquera
XGI Higher-order networks Python Landry et al. (2023)
Visualisation Cytoscape General-purpose networks GUI Shannon et al. (2003)
Gephi General-purpose networks GUI Bastian et al. (2009)
ggraph General-purpose networks R Pedersen (2024)
Graphia General-purpose networks GUI Freeman et al. (2022)
MultiNet Multivariate networks Web-based Wilburn and Lex (2023)
multiNetX Multilayer networks Python Kouvaris
muxViz Multilayer networks R, Web-based De Domenico (2022)
Netminer General-purpose networks GUI Ghim et al. (2014)
Netwulf General-purpose networks Python, Web-based Aslak and Maier (2019)
NodeXL General-purpose networks Excel 251
Pajek General-purpose networks GUI Batagelj and Mrvar (2004)
py3Plex Multilayer networks Python Skrlj et al. (2019)
Sigma.js General-purpose networks JavaScript 302
Tulip General-purpose networks GUI Auber et al. (2012)
Table 2: Libraries and tools for network science analysis and visualisation, with support for multilayer, temporal, and higher-order networks. “Main focus” refers to the area where the software is primarily applied; however, some tools can support multiple network types.
@@ V Outlook
The past decade has shown that multilayer networks are not merely a convenient generalization of graph theory but a fundamentally new modelling language. The field has now reached a stage where the central question is no longer how to adapt single-layer tools, but how to identify the truly multilayer mechanisms that cannot be reduced away. Several theoretical directions emerge from this perspective. One concerns the search for minimal descriptions: determining when layers encode irreducible information and when they can be compressed. Another direction concerns the interplay between multilayer topology and cross-layer dynamical coupling, where instabilities, hybrid phase transitions, metastability, or coexistence of dynamical regimes can arise due to the multiplex nature of the system. Understanding the general principles that govern these phenomena remains a major open challenge. At the same time, the field’s inherently multidisciplinary nature complicates the picture: multilayer approaches are used in physics, biology, economics, computer science, and the social sciences, often with different assumptions, terminologies, and modelling priorities. This diversity makes communication across communities difficult, yet it also underscores the value of multilayer networks as a common representational language capable of unifying disparate viewpoints.
At the methodological frontier, several additional opportunities demand attention. The first involves integrating multilayer representations with temporal and higher-order structures. Many real systems combine heterogeneous interaction types with time-varying group interactions, but current modelling frameworks still tend to treat these features separately. Developing unified theoretical tools capable of capturing polyadic interactions, memory effects, and cross-layer dependencies simultaneously is an important next step. A second direction concerns inference, where the rise of generative models, statistical reconstruction techniques, and machine-learning-based embeddings is beginning to provide new ways to infer hidden layers and mesoscale structures. Yet, understanding the limits of inference remains unresolved. A third challenge lies in controllability and optimisation: while single-layer structural controllability is now well understood, its multilayer analogue is largely unexplored, especially when layers interact through non-linear or time-dependent couplings. Finally, the rapidly growing integration between network science and artificial intelligence opens a promising but underdeveloped path. Deep learning architectures are themselves multilayered systems whose internal dynamics could benefit from tools developed in multilayer network science, potentially contributing to interpretability and robustness in AI.
@@
Looking ahead, progress will depend critically on the availability of high-resolution multilayer data, standardised formats, and interoperable tools, but also on conceptual clarity about what it means for a system to be truly multilayer. If these developments continue in parallel, multilayer network science is poised to move from descriptive analyses to predictive, actionable models capable of guiding interventions, optimising coupled infrastructures, and informing policy in domains where interdependence is no longer optional to ignore. As the complexity of modern systems continues to grow, the multilayer perspective will remain essential for understanding the intertwined processes that characterise the natural and engineered world.
@@ Acknowledgements.
All authors acknowledge the support of the AccelNet-MultiNet program, a project of the National Science Foundation (Award #1927425 and #1927418). We thank M. Clarin from COSNET Lab-BIFI and N. Samay from NetSI for helping with the figures. A.S.T. acknowledges support by FCT – Fundação para a Ciência e Tecnologia – through the LASIGE Research Unit, ref. UID/408/2025. A.A. acknowledges support from the grant RYC2021‐033226‐I funded by MCIN/AEI/10.13039/501100011033 and the European Union NextGenerationEU/PRTR. A.A. and Y.M. were partially supported by the Government of Aragon, Spain, and ERDF ”A way of making Europe” through grant E36-23R (FENOL), and by Grant No. PID2023-149409NB-I00 from Ministerio de Ciencia, Innovación y Universidades, Agencia Española de Investigación (MICIU/AEI/10.13039/501100011033) and ERDF ”A way of making Europe”. G.F.A was partially supported by the São Paulo Research Foundation (FAPESP), through grants 2024/16711-8 and 2025/04409-8. A.D.-G. and O.A. acknowledge support from the Spanish Grant No. PID2021-128005NB-C22, funded by MCIN/AEI/10.13039/501100011033 and “ERDF A way of making Europe;” and from Generalitat de Catalunya (No. 2021SGR00856). M.S. acknowledges support from Grants No. RYC2022-037932-I and CNS2023-144156 funded by MCIN/AEI/10.13039/501100011033 and the European Union NextGenerationEU/PRTR. M.K. acknowledges support from the ANR funded DATAREDUX project (ANR-19-CE46-0008); the National Laboratory for Health Security, Alfréd Rényi Institute of Mathematics, RRF-2.3.1-21-2022-00006; and the WWTF funded MOMA project (10.47379/ESS22032). J.K. gratefully acknowledges partial support by ERC grant No. 810115-DYNASNET. G.P acknowledges support from the European Research Council (ERC) Consolidator Grant under the European Union’s Horizon Europe programme (grant agreement No. 101171380, project RUNES) and the MSCA Doctoral Network BeyondTheEdge(Grant no. 101120085).